OpenStudy (anonymous):

@iGreen 4. -2c=-18 a. c=-9 b. c=-36 c. c=9 d. c=36 5. 2/3x=6 a. x=9 b. x=4 c. x=18 d. x=12 6. 6.5=-3+n a. n=3.5 b. n=9.5 c. n=19.5 d. n=-19.5

OpenStudy (igreen):

-2c = -18 Divide -2 to both sides: -18 / -2 = 9 c = 9

OpenStudy (igreen):

2/3x=6 Divide by 2/3 or multiply by its reciprocal, 3/2. 3/2 * 6 = 9. x = 9

OpenStudy (igreen):

6.5=-3+n Add -3 to both sides 6.5 + -3 = 3.5 n = 3.5